16. Mai 2017: Dienstag

Today’s Activities:

  • Backups:
    • More work on the follow-up questions for the RFI responses.
    • TAPE clones ran to completion in 35 minutes.
    • COMIS group that would be cloned is still running from the weekend.
    • DISK Clones completed at 12:30 (22h 23m).
  • LDAP/Accounts:
    • Three non-authoritative new records:
      • Found all three and set them up to resolve tonight.
    • Stale groups ran, twenty-seven (27) emails sent.
    • Cleaned up two errors in yesterday’s account purge run.
    • Remove entitlements from departing students as requested.

Tags:

Leave a Reply

You must be logged in to post a comment.